OpenStudy (anonymous):

how do you solve dy/dx for y=(x^2 + 3x)(x-2)(x^2 +1) by logarithmic differentiation?

OpenStudy (anonymous):

so ln y = ln(x2+3x) + ln(x-2) + ln(x2 +1) right?

OpenStudy (anonymous):

\[\ln y=\ln (x^2+3x)(x-2)(x^2+1)\]

OpenStudy (anonymous):

\[\ln(abc)=\ln a+\ln b+ \ln c\]

OpenStudy (anonymous):

and we get 1/y (dy/dx) = [(1/ x2+3x)(2x+3)]+ [(1/ x-2)]+ [(1/ x2+1)(2x)]

OpenStudy (anonymous):

im stuck here though...because do i just simplify? or do i have to do anything else?

OpenStudy (anonymous):

dont just leave it as 1 on the nominator but derive the denominator at the top

OpenStudy (anonymous):

\[\frac{dy}{dx}\ln(x^2+3x)=\dfrac{2x+3}{x^2+3x}\]

OpenStudy (anonymous):

right so now we have \[[(2x+3)/(x ^{2} +3x)] +(1/x-2) + (2x/x ^{2} +1)\]

OpenStudy (anonymous):

do i simplify more?

OpenStudy (anonymous):

so multiply by y

OpenStudy (anonymous):

wait why are we doing that?

OpenStudy (anonymous):

since you have \[\frac{\frac{dy}{dx}}{y}\]

OpenStudy (anonymous):

on the left

OpenStudy (anonymous):

so multiply everything by y then?

OpenStudy (anonymous):

\[\frac{dy}{dx}=y[\frac{2x+3}{x^2}+\frac{1}{x-2}+\frac{2x}{x^2+1}]\] \[\frac{2x+3}{x^2+3x}+\frac{1}{x-2}+\frac{2x}{x^2+1}] (x^2+3x)(x-2)(x^2+1)\]

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

so now we have (x2 +3x)(x-2)(x2+1) + (x2+3x)(x2+1) + (2x)(x2+3x)(x-2)

OpenStudy (anonymous):

sorry (2x+3)(x-2)(x2+1) + (x2+3x)(x2+1) + (2x)(x2+3x)(x-2)

OpenStudy (anonymous):

should i leave it like that?

OpenStudy (anonymous):

yes good work

OpenStudy (anonymous):

i dont have to solve the binomials?

OpenStudy (anonymous):

thanks for you help :)

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

yw